OpenStudy (anonymous):

If, the point (2,-1) is on the terminal side of angle 0, find the values of sin0,cos0,tan0

OpenStudy (anonymous):

draw the picture and connect with the origin. you will see a right triangle, so all you need is the length of the hypotenuse. we get that by pythagoras:\[\sqrt{2^2+(-1)^2}=\sqrt{5}\] the coordinates are (2,-1) \[\cos(\theta)=\frac{2}{\sqrt{5}}\] \[\sin(\theta)=\frac{-1}{\sqrt{5}}\] \[tan(\theta)=\frac{-1}{2}\]